Operations Report – March 19th

SOL: 3

Name of person filing report: Jesús, Crew Engineer

Non-nominal systems: None

Notes on non-nominal systems: None

ROVERS

Spirit rover : not used

Opportunity rover: not used

Curiosity rover: Used

Beginning Hours: 302

Beginning charge: 100%

Ending Hours: 302.4

Ending charge: 75%

Currently charging: yes

Perseverance rover: used

Beginning Hours: 311.8

Beginning charge: 100%

Ending Hours: 312

Ending charge: 92%

Currently charging: yes

General notes on rovers: Worked as expected.

Summary of Hab operations: Exercise activities, cooking, EVA preparation, equipment troubleshooting

Water Use (please use both methods to estimate water usage)

Time of measurements: 5:00 pm
1) Per formula (see notes): 22.22 gal
2) Smart Home Dashboard (see notes): 32.83 gal

Water (static tank, remaining gallons): 418.19 gal (based on Smart Home)

Static tank pipe heater (on or off): on

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: used from 7 pm to 10 pm for reports

Summary of suits and radios:

Suits :

1 -> 100%, used

2 -> 100%, used

3 -> 100%, not used

Suits : Bolts on helmet hinges were very loose during/after the first EVA but were tightened for the second.

Radios : Radio 1 stopped working during the last portion of second EVA but resumed functionality after it was turned off and on upon return to HAB. Cause unknown.

Summary of GreenHab operations: Plants were watered.

WATER USE: 6 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 5 to 10 pm

Harvest (name, weight in grams): None

Summary of ScienceDome operations: None

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done): None

Summary of any observatory issues: None

Summary of health and safety issues: None

Sol Summary – March 18th

Crew 309 Sol Summary Report 18Mar2025

Sol: 2

Summary Title:

Author’s name: Co-Commander A

Mission Status: Nominal

Sol Activity Summary:

  • This morning commander A and crew engineer went on their 2nd EVA of the week. Today was especially windy and so the crew was mainly upstairs to keep warm because the heaters were struggling to keep up. Following lunch the crew noticed the smell of propane down stairs and began investigating the heater. We determined that the high winds were affecting the heating system. The high winds also knocked down a safety ladder which the crew had to secure. Eventually we had to shut off the heater for a few hours and just use a small space heater upstairs. As the winds began to slow we were able to turn the heat back on and our second EVA, which was delayed by the weather, was completed successfully. In the evening we had a good dinner and conversed about lessons learned from the day.

Look Ahead Plan: We will be conducting two EVAs tomorrow in the morning and afternoon

Anomalies in work:

  • The heating system was set to 70 and then 72 but the hab stayed consistently at 67. We believe this is due to the strong winds. An hour later it has risen to 68 but has not reached 72. Upon examination, the heating system is not consistently blowing hot air. Current time is about 10:00 AM.
  • Crew smelled a propane smell, and a gas alarm went off. Engineer used a gas detector and no leaks were detected. Crew examined the HVAC system and found an error code blinking. The diagnosis is that the excessive winds are affecting the pressure switch of the HVAC causing it to turn on and off. This diagnosis explains the HVAC not staying continuously on until target temperature is reached and the propane smell. Current temperature is set to reach 72 degrees, but the temperature is at 67. Winds are at about 30 mph and should be less tonight. Current time is about 1:00 PM.
  • Crew communicated with Mission Support to replace the HVAC filter. Crew has replaced the air filter for the HVAC. Current time is about 2:45 PM.
  • The propane smell persisted and became substantial at 3:00 PM. We are turning the HVAC off and using the space heater on the green/low setting.
  • Around 5:30 PM, the wind reduced and the crew turned the HVAC system back on. We have experienced and has had no issues going

Weather: Windy, 27F / 48F

Crew Physical Status: Nominal

EVA: 2 EVAs completed

Reports to be filed: Sol Summary, Journalist Report, GreenHab, Operations Report

Support Requested: None

GreenHab Report – March 18th

GreenHab Officer: Prachi Dutta

Environmental control (fan & heater): Yes

Average temperatures (last 24h): 91.1 F

Maximum temperature (last 24h): 131 F

Minimum temperature (last 24h): 40.8 F

Hours of supplemental light: 5pm-10pm

Daily water usage for crops: 12 Gallons

Daily water usage for research and/or other purposes: None

Water in Blue Tank (200 gallon capacity): 183 gallons

Time(s) of watering for crops: 1.30 pm

Changes to crops: None

Narrative: Watered the crops.

Harvest: 30g basil, 92g of carrots

Support/supplies needed: None

Operations Report – March 18th

SOL: 2

Name of person filing report: Jesús, Crew Engineer

Non-nominal systems: HAB Heating system

Notes on non-nominal systems: The HAB heating system was failing to heat consistently due to excessive winds, replaced air filter, checked for gas leaks, replaced air filter, shut off heating system at around 3 pm due to intense propane smell across the HAB and began use of the electric space heater, but returned to furnace use around 6 pm.

ROVERS

Spirit rover : not used

Opportunity rover: not used

Curiosity rover: Used

Beginning Hours: 301.7

Beginning charge: 100%

Ending Hours: 302

Ending charge: 73%

Currently charging: no

Perseverance rover: used

Beginning Hours: 311.5

Beginning charge: 100%

Ending Hours: 311.8

Ending charge: 75%

Currently charging: yes

General notes on rovers: Worked as expected.

Summary of Hab operations: Heating system troubleshooting, cooking, EVA preparation

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 4:50 pm
1) Per formula (see notes): 45.75 gal
2) Smart Home Dashboard (see notes): 27.35 gal

Water (static tank, remaining gallons): 451.02 gal (based on Smart Home)

Static tank pipe heater (on or off): on

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: used from 7 pm to 10 pm for reports

Summary of suits and radios:

Suits :

1 -> 100%, used

2 -> 100%, used

3 -> 100%, not used

Suits : Nothing to report.

Radios : During EVA 5, chatter between radio 4 and 5 was full of static for around 30 min of the EVA

Summary of GreenHab operations: Plants were watered.

WATER USE: 12 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 5 to 10 pm

Harvest (name, weight in grams):

Basil: 30 g

Carrots: 92 g

Summary of ScienceDome operations: None

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done): Medic worked on medical rover, no specific tools were used. Gas sniffer and space heater were moved from RAM to the main HAB due to heating issues.

Summary of any observatory issues: None

Summary of health and safety issues: None

Journalist Report – March 18th

Journalism Report for Sol 2
Red Mars, Green Mars
Mars is known for its reddish color in the night sky and this crew can confirm that the dirt out here is quite red. Each EVA tracks in some dust and this morning’s crew brought in a whole beach’s worth of sand which we had to sweep up afterwards. The red of the Martian terrain is a unique sort of beautiful, but it’s nothing compared to the wonderful green in our GreenHab. This module might be the best in the whole complex. The GreenHab has lots of tomatoes, onions, and basil which we used to make a tasty spaghetti dinner tonight. Not only is it full of plants and veggies to eat, but it’s always warm – and Mars is very cold. Especially today. The GreenHab was a true oasis of warmth today while our main hab heater struggled to keep up with the wind and cold and our crew engineer tried to fix it.

Sol Summary – March 21st

Crew 309 Sol Summary Report 21Mar2025

Sol: 5

Summary Title: Final Day

Author’s name: Co-Commander A

Mission Status: Nominal

Sol Activity Summary:

  • We conducted 3 EVAs today. GreenHab and Engineer went in the morning, commander A and Medic in the afternoon, and commander B and Scientist last. The crew began the cleaning process completing the RAM, GreenHab, and Food Inventory tasks this evening. Simulation ended after dinner and the crew was able to explore some of the Martian terrain outside of our EVA suits.

Look Ahead Plan: We will be packing up and heading home in the morning.

Anomalies in work:

  • During prep for our second EVA of the day, we determined that the helmet had become detached from the spacesuit worn by crew member one. Engineer, Medic, and Commander A were able to repair the suit for the remainder of the EVAs. A more permanent solution will be required for future missions.

Weather: Windy, 34F / 62F

Crew Physical Status: Nominal

EVA: 3 EVAs completed

Reports to be filed: Sol Summary, Journalist Report, GreenHab, Operations Report, Mission Summary

Support Requested: None

Operations Report – March 21st

SOL: 5

Name of person filing report: Jesús, Crew Engineer

Non-nominal systems: None

Notes on non-nominal systems: None

ROVERS

Spirit rover : used (EVA 2)

Beginning Hours: 268.9

Beginning charge: 100%

Ending Hours: 269.5

Ending charge: 65%

Currently charging: yes

Opportunity rover: used

Beginning Hours: 209.1

Beginning charge: 100%

Ending Hours: 209.5

Ending charge: 73%

Currently charging: yes

Curiosity rover: used

Beginning Hours: 302.5

Beginning charge: 100%

Ending Hours: 302.7

Ending charge: 80%

Currently charging: yes

Perseverance rover: used

Beginning Hours: 312.3

Beginning charge: 100%

Ending Hours: 312.7

Ending charge: 84%

Currently charging: yes

General notes on rovers: Worked as expected.

Summary of Hab operations: Cleaning, cooking, EVA preparation, packing

Water Use (please use both methods to estimate water usage)

Time of measurements: 5:00 pm
1) Per formula (see notes): 33.98 gal
2) Smart Home Dashboard (see notes): 45.34 gal

Water (static tank, remaining gallons): 330.58 gal (based on Smart Home)

Static tank pipe heater (on or off): on

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: will be using from 7 pm onwards for reports, indefinitely due to yesterday’s malfunction

Summary of suits and radios:

Suits :

1 -> 100%, used

2 -> 100%, used

3 -> 100%, not used

Suits : Lost upper-left nut and bolt that fastened helmet to backpack. Replaced using bolt and washers from repair kit but had to use a nut from RAM due to improper fitting of available nuts in repair kit.

Radios : Nothing to report.

Summary of GreenHab operations: Plants were watered.

WATER USE: 9 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 5 to 10 pm

Harvest (name, weight in grams):

Scallions: 69 g

Summary of ScienceDome operations: None

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done): Commandeered hex nut for EVA suit repair.

Summary of any observatory issues: None

Summary of health and safety issues: None

GreenHab Report – March 21st

GreenHab Officer: Prachi Dutta

Environmental control (fan & heater): Yes

Average temperatures (last 24h): 86 F

Maximum temperature (last 24h): 131 F

Minimum temperature (last 24h): 40.8 F

Hours of supplemental light: 5pm-10pm

Daily water usage for crops: 9 Gallons

Daily water usage for research and/or other purposes: None

Water in Blue Tank (200 gallon capacity): 161 gallons

Time(s) of watering for crops: 1.00 pm

Changes to crops: None

Narrative: Watered, cleaned the crops, pulled out the weeds. Sweeped and cleaned the GreenHab for our departure tomorrow.

Harvest: 69g scallions

Support/supplies needed: None

Journalist Report – March 21st

Here is our Journalism report for today.
Strangers to Martians
When we started the Mars mission just a few Sols ago each of us was almost a complete stranger to rest with almost no idea of what we had really signed ourselves up for and now after 5 days in each other’s constant company we have become a true Mars crew. Our crew bonded through the endless training on Sol 0 and then sorting through that same information to determine our crew responsibilities on Sol 1. We got to work with each other to solve problems on EVAs, learning more about each other each time we were paired with someone new. We did daily activities together and exchanged ideas on life habits like how we each like to exercise or how our crew scientist likes to eat random handfuls of spinach to hit her daily veggie count (it’s honestly better than one might expect). We had casual conversations about our lives and our individual research and played games like Pictionary that resulted in an abundance of laughs. We vlogged our Martian cooking, including a few mishaps, and tried lots of dehydrated food for the first time. We gazed up at the stars and wondered whether we could ever explore the entirety of our vast universe. And now on Sol 5 as we approach the end of our mission, we’re conducting our final EVA’s; cleaning the hab; and savoring our last moments on Mars.

Copyright © The Mars Society. All rights reserved. | Main Site